# Operator Spreading in Random Unitary Circuits with Unitary-invariant Gate Distributions

### v2

@@ II.2 Time evolution
$O(t)=U(t,t^{\prime})O(t^{\prime})U(t,t^{\prime})^{\dagger}.$ (9)
@@ II.3 Time evolution of Pauli-string coefficients
$W_{ap}(t,t-1)=\frac{1}{q^{L}}\mbox{tr}\,U(t,t-1)O_{a}U(t,t-1)^{\dagger}O_{p}^{\dagger}.$
@@
${\cal W}_{ap;x,y}=$ $\,\frac{1}{q^{2}}\mbox{tr}\,{\cal U}_{x,y}(\sigma_{a_{x}}\otimes\sigma_{a_{y}}){\cal U}_{x,y}^{\dagger}(\sigma_{p_{x}}^{\dagger}\otimes\sigma_{p_{y}}^{\dagger}).$ (16)
@@
$\sum_{p_{x},p_{y}}{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}=$ $\,\delta_{a_{x},b_{x}}\delta_{a_{y},b_{y}},$
$\sum_{a_{x},a_{y}}{\cal W}_{ap;x,y}{\cal W}_{aq;x,y}=$ $\,\delta_{p_{x},q_{y}}\delta_{q_{x},q_{y}},$ (17)
@@ III.1 Stochastic growth model
where $A_{1}$ is a complex coefficient and $A_{2}$ and $A_{3}$ are real coefficients that are linearly independent combinations of the moment functions,
$A_{1}=$ $\,\frac{2({\cal R}_{2;2}+{\cal R}_{1,1;1,1}-4{\cal R}_{1,1})}{q^{4}(q^{4}-9)}$
@@ III.4 Right-propagating $n$-point density
$\,\mbox{}+\sum_{\bar{a}_{1},\ldots,\bar{a}_{n+1}}T_{\bar{a}_{n+1}\bar{a}_{n};\bar{p}_{n}\bar{p}_{n-1}}\ldots T_{\bar{a}_{3},\bar{a}_{2};\bar{p}_{2},\bar{p}_{1}}T_{\bar{a}_{1}\mathbb{X};\mathbb{X}\mathbb{I}\,}\ldots\bar{\rho}^{(n+1)}(\Delta x+2;t-1;\bar{a}_{n+1},\bar{a}_{n},\ldots,\bar{a}_{1})$
@@ IV Butterfly velocity for Poisson kernel
As an application, we now consider a random unitary circuit for which the two-qudit gate operators ${\cal U}_{x,x+1}$ are taken from the Poisson-kernel distribution (6) with ensemble average $𝟙\langle{\cal U}_{x,x+1}\rangle=\alpha\openone$. The Poisson kernel interpolates between the Haar distribution for $\alpha=0$ and a trivial circuit with no information spreading for $|\alpha|=1$.
Figure 4: Butterfly velocity for a random unitary circuit with Poisson kernel-distributed two-qudit gates for $q=2$ and $q\to\infty$. The interpolation parameter $\alpha$ determines the ensemble average $𝟙\langle{\cal U}_{x,x+1}\rangle=\alpha\openone$ of the two-qudit gate operator ${\cal U}_{x,x+1}$. The solid curves show the approximations $v_{\rm B}^{(n)}$ for order $n=0$, $2$, and $4$. The inset illustrates the differences between the fourth-order and second-order approximations, as well as the differences between the second-order and zeroth-order approximations. For $q\to\infty$ the approximation scheme is exact already for $n=0$. The data points are from a direct numerical simulation of the stochastic process (38) for $L=320$, which is based on $N=2\times 10^{5}$ independent realizations.
Figure 5: Same as Fig. 4, but for the diffusion constant ${\cal D}$ for a random unitary circuit with Poisson kernel-distributed two-qudit gates. The small fluctuations of the data points around the theoretical curve are attributed to the remaining statistical fluctuations of the large, but finite number of realizations $N$ of the stochastic growth process in the numerical simulations.
For the Poisson kernel, the coefficients $A_{1}$, $A_{2}$, and $A_{3}$ in Eq. (28) are (see App. E)
@@
Figures 4 and 5 show these zeroth-order results for butterfly velocity vs. $|\alpha|$ for $q=2$ and for the limit $q\to\infty$, together with a numerical evaluation of the $n$th order approximations $v_{\rm B}^{(n)}$ and ${\cal D}^{(n)}$ for $n=2$ and $n=4$. The data points in Figs. 4 and 5 show the result of a direct numerical simulation of the binary stochastic evolution described by Eq. (38). As depicted in Fig 4, already the zeroth-order approximation is extremely accurate for the butterfly velocity. For the diffusion constant, there is a slightly larger difference between $n=0$ and $n=2$, but here, too, the difference between the approximations at orders $n=2$ and $n=4$ is no longer visible. Hence, for both butterfly velocity and diffusion constant, we conclude that our approximation scheme converges within a distance of only a few sites from the right end of the Pauli string, which suggests that the width $n_{\rm DW}$ of the domain wall between the trivial and maximum-entropy regions of the Pauli string distribution is of that same order.
To further verify the validity of our approximation scheme, we compute the difference between the right-eigenvectors $V_{1}^{(n)}$ of the transition matrix $D^{(n)}+D^{\prime(n)}$ relevant for the information spreading at large times and its approximation according to the Ansatz (III.4). Hereto, we define
@@
In Fig. 6 we show the two-norm $\|V_{1}^{(n)}-{V}_{1}^{(n)\infty}\|_{2}$ vs. the truncation order $n$. The figure shows that the norm exhibits rapid convergence, particularly when the ensemble closely approximates the Haar measure.
@@ Acknowledgements.
We thank Adam Nahum, Guoyi Zhu, Vatasl Dwivedi, and Adam Chaou for their valuable discussions. Special thanks to Yiping Deng for suggestions on figure editing. Financial support was provided by the Einstein Research Unit on Quantum Devices at Freie Universität Berlin.
@@ Appendix A Generalized Pauli matrices
for arbitrary $q\times q$ matrices $A$ and $B$. Because of Eq. (85), the generalized Pauli matrices $\sigma_{k,j}$ are also said to be a unitary $1$-design [18].
@@ Appendix B Calculation of $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}^{*}\rangle$
${\cal W}_{ap;x,y}=$ $\,\frac{1}{q^{2}}\mbox{tr}\,{\cal U}_{x,y}\Sigma_{a}{\cal U}_{x,y}^{\dagger}\Sigma_{p}^{\dagger},$ (93)
@@
with ${\cal V}_{x,x+1}$ a Haar-distributed random unitary matrix that is statistically independent of ${\cal U}_{x,x+1}$, and then average over ${\cal V}_{x,x+1}$. The result of such an average expresses $\langle{\cal W}_{ap;x,x+1}{\cal W}_{bq;x,x+1}\rangle$ in terms of the four moments of the distribution of the two-qudit evolution matrix ${\cal U}_{x,x+1}$ defined in Eq. (27) of the main text. We use the diagrammatic method developed by Beenakker and one of the authors [21] to perform the average over $\equiv{\cal V}_{x,x+1}$. Following Ref. [21], the elements of the Haar-distributed matrix ${\cal V}_{x,x+1}$ are represented by double lines,
@@
For the calculation of the traces over products of up to two generalized Pauli matrices, we now specialize to the case $p=q$, which is the case of interest for the calculation in the main text. Traces containing one or two generalized Pauli matrices can be calculated with the help of the orthonormality relation (86), which gives $\operatorname{tr}\Sigma_{a}=q^{2}\delta_{a,0}$, $\operatorname{tr}\Sigma_{a}\Sigma_{b}^{\dagger}=q^{2}\delta_{a,b}$, $\operatorname{tr}\Sigma_{a}\Sigma_{p}=q^{2}\delta_{a,-p}$, etc. For a trace with a product of three generalized Pauli matrices, we similarly have $\operatorname{tr}\Sigma_{a}\Sigma_{b}^{\dagger}\Sigma_{p}^{\dagger}\operatorname{tr}\Sigma_{p}=q\delta_{p,0}\operatorname{tr}\Sigma_{a}\Sigma_{b}^{\dagger}=q^{2}\delta_{p,0}\delta_{a,b}$ and $\operatorname{tr}\Sigma_{a}\Sigma_{p}^{\dagger}\Sigma_{p}\operatorname{tr}\Sigma_{b}^{\dagger}=q\delta_{b,0}\operatorname{tr}\Sigma_{a}=q^{2}\delta_{a,0}\delta_{b,0}$, etc. Finally, a trace involving four generalized Pauli matrices can be evaluated using the commutation relation (87), e.g., $\operatorname{tr}\Sigma_{a}\Sigma_{p}\Sigma_{b}^{\dagger}\Sigma_{p}^{\dagger}=-\varphi(a,p)\operatorname{tr}\Sigma_{p}\Sigma_{a}\Sigma_{b}^{\dagger}\Sigma_{p}^{\dagger}=-\varphi(a,p)q^{2}\delta_{a,b}$. After evaluating all traces in this manner, all terms contributing to $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}\rangle$ either contain the Kronecker delta symbol $\delta_{a,b}$ or the product $\delta_{a,0}\delta_{b,0}$. In either case, $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}\rangle=0$ if $a\neq b$, which proves Eq. (22) of the main text.
@@
$A_{1}=$ $\,-2q^{2}(q^{4}-1)B_{3},$
$A_{2}=$ $\,q^{4}(q^{4}-1)B_{5},$
$A_{3}=$ $\,q^{4}(q^{4}-1)B_{6}.$ (104)
The coefficient $A_{1}$ is complex because $B_{3}$ is complex, whereas $A_{2}$ and $A_{3}$ are real.
@@ Appendix D Drift-diffusion equation
Appendix E Poisson kernel

### v3

@@ II.2 Time evolution
$O(t)=U^{\dagger}(t,t^{\prime})O(t^{\prime})U(t,t^{\prime}).$ (9)
@@ II.3 Time evolution of Pauli-string coefficients
$W_{ap}(t,t-1)=\frac{1}{q^{L}}\mbox{tr}\,U^{\dagger}(t,t-1)O_{a}U(t,t-1)O_{p}^{\dagger}.$
@@
${\cal W}_{ap;x,y}=$ $\,\frac{1}{q^{2}}\mbox{tr}\,{\cal U}^{\dagger}_{x,y}(\sigma_{a_{x}}\otimes\sigma_{a_{y}}){\cal U}_{x,y}(\sigma_{p_{x}}^{\dagger}\otimes\sigma_{p_{y}}^{\dagger}).$ (16)
@@
$\sum_{p_{x},p_{y}}{\cal W}_{ap;x,y}{\cal W}^{*}_{bp;x,y}=$ $\,\delta_{a_{x},b_{x}}\delta_{a_{y},b_{y}},$
$\sum_{a_{x},a_{y}}{\cal W}_{ap;x,y}{\cal W}^{*}_{aq;x,y}=$ $\,\delta_{p_{x},q_{y}}\delta_{q_{x},q_{y}},$ (17)
@@ III.1 Stochastic growth model
where $A_{1}$ is a complex coefficient and $A_{2}$ and $A_{3}$ are real coefficients that are linearly independent combinations of the moment functions (for details, we refer to APP. B).
@@ III.4 Right-propagating $n$-point density
$\,\mbox{}+\sum_{\bar{a}_{1},\ldots,\bar{a}_{n+1}}T_{\bar{a}_{n+1}\bar{a}_{n};\bar{p}_{n}\bar{p}_{n-1}}\ldots T_{\bar{a}_{3},\bar{a}_{2};\bar{p}_{2},\bar{p}_{1}}T_{\bar{a}_{1}\mathbb{X};\mathbb{X}\mathbb{I}\,}\bar{\rho}^{(n+1)}(\Delta x+2;t-1;\bar{a}_{n+1},\bar{a}_{n},\ldots,\bar{a}_{1})$
@@ IV Butterfly velocity for Poisson kernel
As an application, we now consider a random unitary circuit for which the two-qudit gate operators ${\cal U}_{x,x+1}$ are taken from the Poisson-kernel distribution (6) with ensemble average $𝟙\langle{\cal U}_{x,x+1}\rangle=\alpha\openone$. The Poisson kernel interpolates between the Haar distribution for $\alpha=0$ and a trivial circuit with no information spreading for $|\alpha|=1$. An alternative interpolating random matrix ensemble for the two-qudit evolution operators can be obtained from the Dyson’s Brownian motion ensemble of random matrices [20, 21, 22]. This interpolating ensemble is discussed in App. E.
Figure 4: Butterfly velocity for a random unitary circuit with Poisson kernel-distributed two-qudit gates for $q=2$ and $q\to\infty$. The interpolation parameter $\alpha$ determines the ensemble average $𝟙\langle{\cal U}_{x,x+1}\rangle=\alpha\openone$ of the two-qudit gate operator ${\cal U}_{x,x+1}$. The solid curves show the approximations $v_{\rm B}^{(n)}$ for order $n=0$, $2$, and $4$. The inset illustrates the relative differences between the fourth-order and second-order approximations, as well as the relative differences between the second-order and zeroth-order approximations. For $q\to\infty$ the approximation scheme is exact already for $n=0$. The data points are from a direct numerical simulation of the stochastic process (35) for $L=320$, which is based on $N=2\times 10^{5}$ independent realizations. Due to the rapid convergence of the truncation scheme, the analytical results for orders $n=0$, $2$, and $4$ and the numerical simulation data in the top panel are indistinguishable.
Figure 5: Same as Fig. 4, but for the diffusion constant ${\cal D}$ for a random unitary circuit with Poisson kernel-distributed two-qudit gates. The small fluctuations of the data points around the theoretical curve are attributed to the remaining statistical fluctuations of the large, but finite number of realizations $N$ of the stochastic growth process in the numerical simulations. Due to the rapid convergence of the truncation scheme, the analytical results for orders $n=$ $2$ and $4$ in the top panel are indistinguishable.
For the Poisson kernel, the coefficients $A_{1}$, $A_{2}$, and $A_{3}$ in Eq. (28) are (see App. F)
@@
For $|\alpha|$ close to $1$, the probability that the front moves in a single time step becomes small. In this limit, the discrete time evolution of the random circuit can be approximated by a continuous random process. Such a continuous random process is again characterized by a drift velocity and diffusion constant [10, 12]. To find these within our approach, we set $|\alpha|^{2}=e^{-\lambda}$, rescale $t^{\prime}=\lambda t$, and take the limit $\lambda\to 0$. Using $t^{\prime}$ as the time variable for the random continuous process, the lowest-order approximations for the drift velocity and diffusion constant become
$v_{\rm B}^{\prime(0)}=$ $\,\frac{(q^{2}-1)^{2}(q^{2}+2)}{q^{2}(q^{2}+1)(q^{2}+3)},$ (65)
${\cal D}^{\prime(0)}=$ $\,\frac{2(q^{4}+q^{2}-2)}{q^{2}(q^{2}+3)}.$ (66)
Figures 4 and 5 show the zeroth-order results for the butterfly velocity vs. $|\alpha|$ for $q=2$ and for the limit $q\to\infty$, together with a numerical evaluation of the $n$th order approximations $v_{\rm B}^{(n)}$ and ${\cal D}^{(n)}$ for $n=2$ and $n=4$. The data points in Figs. 4 and 5 show the result of a direct numerical simulation of the binary stochastic evolution described by Eq. (35). As depicted in Fig 4, already the zeroth-order approximation is extremely accurate for the butterfly velocity. This includes the limit $|\alpha|\to 1$, in which the butterfly velocity goes to zero and the random circuit becomes a random continuous process. For the diffusion constant, there is a slightly larger difference between $n=0$ and $n=2$, but here, too, the difference between the approximations at orders $n=2$ and $n=4$ is no longer visible. Hence, for both butterfly velocity and diffusion constant, we conclude that our approximation scheme converges within a distance of only a few sites from the right end of the Pauli string, which suggests that the width $n_{\rm DW}$ of the domain wall between the trivial and maximum-entropy regions of the Pauli string distribution is of that same order.
Sufficiently far from the moving front, the maximally random approximation of Eq. (III.4) should hold for the $n$-point density $\rho^{(n)}(\Delta x;t;\bar{a}_{n},\ldots,\bar{a}_{1})$. In the long-time limit, the information on the dependence of the $n$-point density on the Pauli string elements $\bar{a}_{n}$, …, $\bar{a}_{1}$ is encoded in the right-eigenvector $V_{1}^{(n)}$ of the transition matrix $D^{(n)}+D^{\prime(n)}$, see Eq. (50). To further verify the validity of our approximation scheme, we compute the difference between the right-eigenvectors $V_{1}^{(n)}$ of the transition matrix $D^{(n)}+D^{\prime(n)}$ relevant for the information spreading at large times and its approximation according to the Ansatz (III.4). Hereto, we define
@@
(We compare $V^{(n)}$ and $V^{(n-2)}$ to avoid even-odd effects intrinsic to the brickwork structure of the random circuit.) In Fig. 6 we show the two-norm $\|V_{1}^{(n)}-{V}_{1}^{(n)\infty}\|_{2}$ vs. the truncation order $n$. The figure shows that the norm exhibits rapid convergence, particularly when the ensemble closely approximates the Haar measure.
@@ Acknowledgements.
We thank Adam Nahum, Guoyi Zhu, Vatasl Dwivedi, and Adam Chaou for their valuable discussions. Special thanks to Yiping Deng for suggestions on figure editing. Financial support was provided by the Einstein Stiftung Berlin (Einstein Research Unit on Quantum Devices).
@@ Appendix A Generalized Pauli matrices
for arbitrary $q\times q$ matrices $A$ and $B$. Because of Eq. (84), the generalized Pauli matrices $\sigma_{a}$ are also said to be a unitary $1$-design [18].
@@ Appendix B Calculation of $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}^{*}\rangle$
${\cal W}_{ap;x,y}=$ $\,\frac{1}{q^{2}}\mbox{tr}\,{\cal U}^{\dagger}_{x,y}\Sigma_{a}{\cal U}_{x,y}\Sigma_{p}^{\dagger},$ (92)
@@
with ${\cal V}_{x,x+1}$ a unitary matrix that can be chosen arbitrarily. In this appendix, we first present a simple argument in support of Eq. (22), making a specific choice for the matrices ${\cal V}_{x,x+1}$. After that, we take the matrices ${\cal V}_{x,x+1}$ to be Haar-distributed random matrices that are statistically independent of ${\cal U}_{x,x+1}$ and perform an average over ${\cal V}_{x,x+1}$. This allows us to derive Eq. 28), which expresses the ensemble average $\langle{\cal W}_{ap}{\cal W}_{bp}^{*}\rangle$ in terms of the four moments of the pair-qudit evolution operator ${\cal U}_{x,x+1}$ defined in Eq. (27) of the main text.
Proof of Eq. (22).— After the substitution (93), we can rewrite the average $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}^{*}\rangle$ as
$\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}^{*}\rangle$ $=\frac{1}{q^{4}}\mbox{tr}\,{\cal U}^{\dagger}_{x,y}\widetilde{\Sigma}_{a}{\cal U}_{x,y}\widetilde{\Sigma}_{p}^{\dagger}\mbox{tr}\,{\cal U}^{\dagger}_{x,y}\widetilde{\Sigma}^{\dagger}_{b}{\cal U}_{x,y}\widetilde{\Sigma}_{p}$ (94)
where
$\widetilde{\Sigma}_{a}=\mathcal{V}_{x,y}\Sigma_{a}\mathcal{V}^{\dagger}_{x,y}.$ (95)
We make the special choice that $\mathcal{V}_{x,y}=\Sigma_{q}$ is a generalized Pauli matrix. From Eq. (86), we then find, that
$\widetilde{\Sigma}_{a}=-\varphi^{*}(a,q)\Sigma_{a},$ (96)
with $\varphi(a,q)$ defined in Eq. (87). It follows that
$\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}^{*}\rangle=-\varphi(b-a,q)\langle{\cal W}_{ap}{\cal W}_{bp}^{*}\rangle,$ (97)
where we made use of the property $\varphi(a,q)^{*}\varphi(b,q)=-\varphi(b-a,q)$, which immediately follows from the definition (87). Because $\Sigma_{q}$ can be chosen arbitrary, we conclude that $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}^{*}\rangle$ can be nonzero only if $\Sigma_{a}=\Sigma_{b}$, which immediately leads to Eq. (22).
Proof of Eq. (28).— To derive Eq. (28), we use the diagrammatic method developed by Beenakker and one of the authors [24] to perform the average over $\equiv{\cal V}_{x,x+1}$. Following Ref. [24], the elements of the Haar-distributed matrix ${\cal V}_{x,x+1}$ are represented by double lines,
@@
For the calculation of the traces over products of up to two generalized Pauli matrices, we now specialize to the case $p=q$, which is the case of interest for the calculation in the main text. Traces containing one or two generalized Pauli matrices can be calculated with the help of the orthonormality relation (85), which gives $\operatorname{tr}\Sigma_{a}=q^{2}\delta_{a,0}$, $\operatorname{tr}\Sigma_{a}\Sigma_{b}^{\dagger}=q^{2}\delta_{a,b}$, $\operatorname{tr}\Sigma_{a}\Sigma_{p}=q^{2}\delta_{a,-p}$, etc. For a trace with a product of three generalized Pauli matrices, we similarly have $\operatorname{tr}\Sigma_{a}\Sigma_{b}^{\dagger}\Sigma_{p}^{\dagger}\operatorname{tr}\Sigma_{p}=q\delta_{p,0}\operatorname{tr}\Sigma_{a}\Sigma_{b}^{\dagger}=q^{2}\delta_{p,0}\delta_{a,b}$ and $\operatorname{tr}\Sigma_{a}\Sigma_{p}^{\dagger}\Sigma_{p}\operatorname{tr}\Sigma_{b}^{\dagger}=q\delta_{b,0}\operatorname{tr}\Sigma_{a}=q^{2}\delta_{a,0}\delta_{b,0}$, etc. Finally, a trace involving four generalized Pauli matrices can be evaluated using the commutation relation (86), e.g., $\operatorname{tr}\Sigma_{a}\Sigma_{p}\Sigma_{b}^{\dagger}\Sigma_{p}^{\dagger}=-\varphi(a,p)\operatorname{tr}\Sigma_{p}\Sigma_{a}\Sigma_{b}^{\dagger}\Sigma_{p}^{\dagger}=-\varphi(a,p)q^{2}\delta_{a,b}$. After evaluating all traces in this manner, all terms contributing to $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}\rangle$ either contain the Kronecker delta symbol $\delta_{a,b}$ or the product $\delta_{a,0}\delta_{b,0}$. In either case, $\langle{\cal W}_{ap;x,y}{\cal W}_{bp;x,y}\rangle=0$ if $a\neq b$, which provides an alternative proof of Eq. (22) of the main text.
@@
$A_{1}=$ $\,\frac{2({\cal R}_{2;2}+{\cal R}_{1,1;1,1}-4{\cal R}_{1,1})}{q^{4}(q^{4}-9)}$
The coefficient $A_{1}$ is complex, whereas $A_{2}$ and $A_{3}$ are real.
@@ Appendix D Drift-diffusion equation
Appendix E Brownian-motion ensemble
The Poisson kernel offers a continuous interpolation between a random circuit with Haar-distributed two-qudit evolution operators and the trivial circuit. Another random matrix ensemble that interpolates between these two limits is Dyson’s Brownian motion ensemble [20, 21]. It obtains the two-qudit operators ${\cal U}_{x,x+1}$ from the continuous time evolution with a random Hamiltonian with Gaussian white-noise distribution. Specifically,
[…]
where $T_{t}$ indicates the time-ordering and ${\cal H}(t^{\prime})$ is a $q^{2}\times q^{2}$ hermitian matrix with Gaussian distribution with zero mean and variance
$\langle H_{ij}(t)H_{kl}(t^{\prime})\rangle=\frac{\lambda}{q^{2}}\delta_{il}\delta_{jk}\delta(t-t^{\prime}),$ (121)
where $\lambda$ is a parameter that describes the interpolation between the trivial and Haar-random two-qudit gate operators. This ensemble satisfies the unitary invariance property (5). The distribution of the eigenphases $e^{i\theta_{j}}$ of ${\cal U}$, $j=1,2,\ldots,q^{2}$, is [30]
$P(\theta_{1},\ldots,\theta_{q^{2}})=$ $\,\frac{2^{\frac{1}{2}q^{2}(q^{2}-1)}}{(q^{2})!}e^{\frac{\lambda}{24}(q^{4}-1)}\prod_{j<k}\sin\frac{\theta_{j}-\theta_{k}}{2}$
×$\,\mbox{}\times\det\left[\frac{f^{(j)}(\theta_{i})}{j!}\right]_{i,j=1,\ldots,q^{2}},$ (122)
where
$f(\theta)=$ $\,\frac{1}{2\pi}\sum_{k=-\infty}^{\infty}e^{-k^{2}\lambda/2q^{2}+ik\theta}$
$=$ $\,\sqrt{\frac{q^{2}}{2\pi\lambda}}\sum_{k=-\infty}^{\infty}e^{-q^{2}(\theta-2\pi k)^{2}/2\lambda}$ (123)
and $f^{(j)}$ denotes the $j$th derivative.
The moments ${\cal R}_{1,1}$, ${\cal R}_{1,1;1,1}$, ${\cal R}_{2;2}$, and ${\cal R}_{1,1;2}$ are [22]
$\mathcal{R}_{1;1}=$ $\,q^{4}e^{-\lambda}+1-e^{-\lambda},$ (124)
$\mathcal{R}_{2;2}=$  $\,2+\frac{1}{2}e^{-2\lambda}\left(q^{4}(q^{4}-3)\cosh\frac{2\lambda}{q^{2}}\right.$
$\,\left.\ \ \ \ \mbox{}-2q^{6}\sinh\frac{2\lambda}{q^{2}}-q^{8}+5q^{4}-4\right),$ (125)
$\mathcal{R}_{1,1;2}=$ $\,e^{-2\lambda}q^{4}\left[q^{2}\cosh\frac{2\lambda}{q^{2}}+\sinh\frac{2\lambda}{q^{2}}\right],$ (126)
$\mathcal{R}_{1,1;1,1}=$ $\,{\cal R}_{2;2}+4(q^{4}-1)e^{-\lambda}+(q^{8}-5q^{4}+4)e^{-2\lambda}.$ (127)
From Eqs. (III.3), (55), (56), and (B)–(109), we then find that the lowest-order approximations for the butterfly velocity and the diffusion constant are
$v_{\rm B}^{(0)}=$ $\,\frac{q^{2}-1}{q^{2}+1}\frac{c-d}{c+d}$ (128)
${\cal D}^{(0)}=$ $\,\frac{4}{(q^{2}+1)^{2}}\frac{(c-d)(c+q^{2}d)(q^{2}c+d)}{c(c+d)^{2}},$ (129)
where
$c=$ $\,e^{2\lambda}(q^{6}-9q^{2})$ (130)
$d=$ $\,(q^{6}-9q^{2})\cosh(\frac{2\lambda}{q^{2}})-2(q^{4}+3)\sinh(\frac{2\lambda}{q^{2}}).$ (131)
In the limit $q\to\infty$ these estimates agree with what we find for the Poisson kernel distribution, provided we identify $|\alpha|^{2}=e^{-\lambda}$.
The Brownian motion ensemble can also be used to construct a continuous random process. To find the drift velocity and the diffusion constant within our approach, we rescale $t^{\prime}=\lambda t$ and take the limit $\lambda\to 0$. Using $t^{\prime}$ as the time variable for the random continuous process, the lowest-order approximations for the drift velocity and diffusion constant become
$v_{\rm B}^{\prime(0)}=$ $\,\frac{(q^{2}-1)^{2}(q^{4}-6)}{q^{4}(q^{4}-9)},$ (132)
${\cal D}^{\prime(0)}=$ $\,\frac{2(q^{4}-1)(q^{4}-6)}{q^{4}(q^{4}-9)}.$ (133)
For large $q$, these expressions agree with the results obtained for the Poisson kernel, Eqs. (65) and (66).
Appendix F Poisson kernel